OpenStudy (anonymous):

please help!!!! i need to know the horizontal tangents of sin(25x) and the vertical tangents of cos(25x). I got answers of 0 and pi/25 and 25pi/2 and 75pi/2, but my online hw says that theyre wrong

OpenStudy (richyw):

\(f(x)=\cos(25x)\) has no vertical tangent lines, are you in cartesian coordinates?

OpenStudy (richyw):

for the first one you can just take the derivative and set it equal to zero. so \[\frac{d}{dx}(\sin25x)=\cos (25x)\cdot\frac{d}{dx}25x=25\cos (25x)\]

OpenStudy (anonymous):

yeah the interval is from 0 to 2pi

OpenStudy (anonymous):

ahh im confused

OpenStudy (anonymous):

i thought that cos(25x), when set equal to 0 yielded pi/50 and 3pi/50